OpenStudy (anonymous):

uppose that for a company manufacturing calculators, the cost, and revenue equations are given by C=90000+40x R=300−x^2/20 where the production output in one week is x calculators. If the production rate is increasing at a rate of 500 calculators when the production output is 6000 calculators, find each of the following: Rate of change in cost = Rate of change in revenue = Rate of change in profit =

OpenStudy (anonymous):

rate of change in cost = C' = 40 * x' = 40 * 500 = ? rate of change of revenue = R' = \(\Large{-2x\times x'\over20}=-{6000\times400\over10}=-\) i.e., decreasing Profit = Revenue - cost \[P(x)=R(x)-C(x)\implies P'(x)=R'(x)-C'(x)=?\]

OpenStudy (anonymous):

so for the cost i got 20000 but for the revenue i calculated the equation and it tells me its the wrong answer

OpenStudy (anonymous):

what value did you put for revenue rate?

OpenStudy (anonymous):

for the revenue i got 240000

OpenStudy (anonymous):

what about the sign?

OpenStudy (anonymous):

I made a note on the sign up there.....

OpenStudy (anonymous):

for the sign was negative

OpenStudy (anonymous):

and the negative answer is wrong?

OpenStudy (anonymous):

@91004775 ok. I'll be turning in now!

OpenStudy (anonymous):

yeah i did tell me it was wrong but thanks anyways